# Conditional Latent Space Molecular Scaffold Optimization for Accelerated Molecular Design

### v1

@@ Abstract
The rapid discovery of new chemical compounds is essential for advancing global health and developing treatments. While generative models show promise in creating novel molecules, challenges remain in ensuring the real-world applicability of these molecules and finding such molecules efficiently. To address this, we introduce Conditional Latent Space Molecular Scaffold Optimization (CLaSMO), which combines a Conditional Variational Autoencoder (CVAE) with Latent Space Bayesian Optimization (LSBO) to modify molecules strategically while maintaining similarity to the original input. Our LSBO setting improves the sample-efficiency of our optimization, and our modification approach helps us to obtain molecules with higher chances of real-world applicability. CLaSMO explores substructures of molecules in a sample-efficient manner by performing BO in the latent space of a CVAE conditioned on the atomic environment of the molecule to be optimized. Our experiments demonstrate that CLaSMO efficiently enhances target properties with minimal substructure modifications, achieving state-of-the-art results with a smaller model and dataset compared to existing methods. We also provide an open-source web application that enables chemical experts to apply CLaSMO in a Human-in-the-Loop setting.
@@ 1 Introduction
Generative models offer various approaches to exploring and creating new chemical compounds. A common strategy involves training a generative model on a comprehensive database of chemical compounds. Once trained, the model can generate entirely new compounds (Gómez-Bombarelli et al., 2018, Tripp et al., 2020, Griffiths & Hernández-Lobato, 2020, Boyar & Takeuchi, 2024, Grosnit et al., 2021, Boyar et al., 2024). This approach opens the door to discovering novel molecules that are unlike any currently known, potentially revealing a vast, untapped universe of chemical possibilities. However, while the generation of these novel molecules is exciting, their practical applicability is often constrained. The challenges lie in synthesizing these novel molecules and in the limited understanding of their properties, which makes it difficult for domain experts to assess their viability Lim et al. (2020).
@@
To address these challenges, another strategy for designing new molecules focuses on modifying/editing existing compounds using generative models Bradshaw et al. (2019), Lim et al. (2020), reinforcement learning Gottipati et al. (2020), genetic algorithms Jensen (2019), or from the domain experts themselves by trial and error Bemis & Murcko (1996), Schreiber (2000), Welsch et al. (2010). These approaches are more likely to produce synthesizable molecules because the base molecule is known to exist and, therefore, can be more tractable for real-world applications. However, even with this strategy, a significant obstacle remains: sample efficiency, i.e., how efficiently a method finds a promising molecular modification with a limited number of molecular property evaluations. Evaluating the properties of a chemical compound is a time-consuming and costly process, and many existing methods in the literature require numerous trials, making them less practical for accelerated discovery. Consequently, there is a critical need for a methodology that can generate target chemical compounds in a more sample-efficient manner.
@@
In this study, we introduce Conditional Latent Space Molecular Scaffold Optimization (CLaSMO) method, a framework designed to address two critical challenges in chemical compound discovery: real-world applicability and sample efficiency. CLaSMO combines a Conditional Variational Autoencoder (CVAE) Higgins et al. (2016) with Latent Space Bayesian Optimization (LSBO) Gómez-Bombarelli et al. (2018) to strategically modify input molecules and optimize their chemical properties.
@@
Figure 1: An example of the CLaSMO framework updating a scaffold for QED optimization. CLaSMO identifies optimal regions in the CVAE latent space and selects bonding points on the scaffold. Chemical features from these points are used to guide the generation of substructures, which are then integrated into the scaffold (A, B, C, D) through small, targeted modifications to improve molecular properties.
@@
In drug discovery, a common strategy for generating synthesizable molecules is to work with molecular scaffolds—key substructures that serve as the foundation for chemical modifications and drug design Bemis & Murcko (1996). Building on this approach, our framework integrates small substructures into these scaffolds to improve key molecular properties. For modifications to be both effective and synthesizable, it’s crucial that the generative model understands how new substructures bond with the scaffold and enhance its properties. To achieve this, we condition substructure generation on the scaffold’s chemical features using a CVAE. Our novel data preparation and training strategy enables the CVAE to generate substructures that align with specific atomic environments. We further optimize this process using LSBO, which efficiently explores both the latent space of the CVAE and the scaffold’s chemical features. This allows CLaSMO to effectively select regions for modification and generate substructures that are chemically meaningful. By tracking molecular similarity between the initial scaffold and the optimized molecule, CLaSMO ensures that modifications are realistic and applicable in real-world settings. Ultimately, this approach accelerates the discovery of novel, synthesizable compounds with improved properties, as illustrated in Figure 1.
We evaluate our approach in two scenarios: optimizing the Quantitative Estimate of Drug-likeness (QED) in Section 5.2 and improving docking simulation scores in Section 5.3. QED assesses how likely a compound is to become a viable drug, while docking scores measure how well a molecule binds to a protein target using a simulator Schrödinger (2023). Our experiments demonstrate that CLaSMO efficiently improves both QED and docking scores in a sample-efficient manner, highlighting its effectiveness. Contributions of this study can be listed as follows:
@@
    We demonstrate that CLaSMO achieves state-of-the-art results using a significantly smaller model and training dataset than its competitors, highlighting its efficiency.
  4. 4.
@@ 2 Related Works
Molecular design strategies can broadly be divided into two categories: from-scratch-generation of molecules and modification-based approaches. Both categories have made significant strides in recent years, yet they also face unique challenges, particularly regarding real-world applicability and sample efficiency.
@@
From-scratch-generation approaches focus on creating entirely new molecules by conducting a search to optimize the target property. A seminal work by Gómez-Bombarelli et al. (2018) introduced latent space optimization-based methodology, using a VAE to generate novel compounds by navigating the latent space of molecular representations. LSBO builds on this by efficiently reducing the number of expensive black-box evaluations required for molecular optimization, enabling the discovery of compounds with desirable properties in a continuous latent space. Since then, numerous studies have further refined and expanded the LSBO framework, focusing on method development and practical applications Grosnit et al. (2021), Tripp et al. (2020), Maus et al. (2022), Griffiths & Hernández-Lobato (2020), Boyar & Takeuchi (2024), Boyar et al. (2024). However, like many other generation-from-scratch methods, such methodologies struggle with real-world applicability—i.e., the difficulty of synthesizing the generated molecules in real-world settings Lim et al. (2020). Other generation methods, such as genetic algorithms Jensen (2019), Grammar VAE Kusner et al. (2017), and Junction Tree (JT) VAE Jin et al. (2018), aim to improve the chemical validity of generated structures. Recent advancements like GP-MOLFORMER Ross et al. (2024) utilizes large language model-like approaches for molecular design, but the real-world applicability challenge remains a major limitation across these methodologies.
@@
Modification-based approaches, on the other hand, focus on adding substructures to existing molecules or scaffolds, often leading to more synthesizable and interpretable designs. Methods like Scaffold-GGM Lim et al. (2020) employ a graph generative model to modify molecular scaffolds, thus improving properties with a higher chance of obtaining synthesizable molecules. Weller & Rohs (2024) introduces DrugHIVE, a deep hierarchical variational autoencoder that leverages scaffold modification to generate novel molecular compounds. There are many other scaffold-based optimization methodologies Li et al. (2019), Langevin et al. (2020), not limited to generative modeling Schreiber (2000), Welsch et al. (2010), Miao et al. (2011). Techniques like Bradshaw et al. (2019)’s model ensure chemical validity in molecular modifications, and Gottipati et al. (2020)’s PGFS model uses reinforcement learning to guide additions to the base molecule. These approaches aim to avoid the low real-world viability problem faced by generation-from-scratch methods, as they build upon known molecular scaffolds, however, they lack advanced optimization methodologies that take sample efficiency into account.
@@
CLaSMO combines the strengths of both categories, leveraging LSBO in the latent space of a CVAE for improved sample efficiency while focusing on scaffold-based modifications. Unlike current LSBO-based molecular design methodologies in the literature, CLaSMO does not generate molecules from scratch but instead optimizes molecular properties by adding substructures to existing scaffolds. This approach mitigates the real-world applicability problem, and increases the chance of obtaining molecules that are both effective and practical for real-world synthesis. By incorporating a targeted scaffold-modification strategy, CLaSMO offers a balanced and efficient solution to molecular design.
@@ 3.1 Conditional Variational Autoencoders (CVAEs)
A VAE Kingma & Welling (2014) consists of an encoder $f_{\phi}^{\text{enc}}:\mathcal{X}\to\mathcal{Z}$ and a decoder $f_{\theta}^{\text{dec}}:\mathcal{Z}\to\mathcal{X}$, where $\mathcal{X}$ represents the input space and $\mathcal{Z}$ the latent space. CVAEs extend the framework of VAEs by incorporating additional condition vector $\bm{c}$ into the latent variable model, facilitating the controlled generation of new instances. In the CVAE architecture, the encoder $q_{\phi}(\bm{z}|\bm{x},\bm{c})$ maps an input $\bm{x}$ and a condition $\bm{c}$ to a latent representation $\bm{z}$. Simultaneously, the decoder $p_{\theta}(\bm{x}|\bm{z},\bm{c})$ reconstructs $\bm{x}$ using both $\bm{z}$ and $\bm{c}$. The training of CVAEs is formulated as the minimization of the conditional variational lower bound:
$\mathcal{L}(\theta,\phi;\bm{x},\bm{c})=-\mathbb{E}_{q_{\phi}(\bm{z}|\bm{x},\bm{c})}\left[\log p_{\theta}(\bm{x}|\bm{z},\bm{c})\right]+\mathrm{KL}\left(q_{\phi}(\bm{z}|\bm{x},\bm{c})\parallel p(\bm{z}\right)),$ (1)
@@ 3.2 Latent Space Bayesian Optimization (LSBO)
In BO, we start with numerous unlabeled instances $\{{\bm{x}_{i}}\}_{i\in[\mathcal{U}]}$ and a smaller set of labeled instances ${(\bm{x}_{i},y_{i})}_{i\in[\mathcal{L}]}$, where an input $\bm{x}_{i}\in\mathcal{X}$ represents a chemical compound, and a label $y_{i}\in\mathcal{Y}\subseteq\mathcal{R}$ indicates its properties such as docking scores. BO seeks to optimize a costly black-box function $f^{\rm BB}:\mathcal{X}\to\mathcal{Y}$, which corresponds to obtaining the physical properties of chemical compounds through experiments or time-consuming simulations in the context of molecular design problems. The goal is to maximize $f^{\rm BB}$ with minimal evaluations, using typically a Gaussian Process (GP) surrogate, trained using the labeled instances $\mathcal{L}$, to predict the function over $\mathcal{X}$. BO uses the surrogate to select an input $\bm{x}$ that may yield values surpassing the current maximum $\max_{i\in\mathcal{L}}y_{i}$. However, building a GP surrogate in high-dimensional spaces like chemical compounds is challenging. LSBO tackles this by employing a VAE/CVAE trained on the unlabelled instances $\mathcal{U}$ to reduce dimensionality, encoding instance in $\mathcal{X}$ to lower dimensional latent space $\mathcal{Z}$. This simplifies surrogate modeling and optimization because $\mathcal{Z}$ has lower-dimension than $\mathcal{X}$. During LSBO iterations, the acquisition function applied to the GP’s predictions selects new points in $\mathcal{Z}$ to evaluate. The chosen latent variable $\bm{z}_{i\textquoteright}$ is decoded into a new input $\bm{x}_{i\textquoteright}=f_{\theta}^{\rm dec}(\bm{z}_{i^{\prime}})$. This new instance is evaluated by $f^{\rm BB}$, and the results update $\mathcal{L}$ and refine the GP model. This cycle repeats until optimal results are achieved or resources are exhausted. In contexts like molecular design, LSBO aims to discover chemical compounds with optimal properties by efficiently navigating the reduced latent space.

### v2

@@ Abstract
The rapid discovery of new chemical compounds is essential for advancing global health and developing treatments. While generative models show promise in creating novel molecules, challenges remain in ensuring the real-world applicability of these molecules and finding such molecules efficiently. To address this challenge, we introduce Conditional Latent Space Molecular Scaffold Optimization (CLaSMO), which integrates a Conditional Variational Autoencoder (CVAE) with Latent Space Bayesian Optimization (LSBO) to strategically modify molecules while preserving similarity to the original input, effectively framing the task as constrained optimization. Our LSBO setting improves the sample-efficiency of the molecular optimization, and our modification approach helps us to obtain molecules with higher chances of real-world applicability. CLaSMO explores substructures of molecules in a sample-efficient manner by performing BO in the latent space of a CVAE conditioned on the atomic environment of the molecule to be optimized. Our extensive evaluations across diverse optimization tasks—including rediscovery, docking score, and multi‑property optimization—show that CLaSMO efficiently enhances target properties, delivers remarkable sample-efficiency crucial for resource‑limited applications while considering molecular similarity constraints, achieves state of the art performance, and maintains practical synthetic accessibility. We also provide an open-source web application [^1] The source code of this work is available at https://github.com/onurboyar/CLASMO-TMLR. that enables chemical experts to apply CLaSMO in a Human-in-the-Loop setting.
@@ 1 Introduction
Generative models offer various approaches to exploring and creating new chemical compounds. A common strategy involves training a generative model on a comprehensive database of chemical compounds. Once trained, the model can generate entirely new compounds (Gómez-Bombarelli et al., 2018, Tripp et al., 2020, Griffiths & Hernández-Lobato, 2020, Boyar & Takeuchi, 2024, Grosnit et al., 2021, Boyar et al., 2024). This approach opens the door to discovering novel molecules that are unlike any currently known, potentially revealing a vast, untapped universe of chemical possibilities. However, while the generation of these novel molecules is exciting, their practical applicability is often constrained. The challenges lie in synthesizing these novel molecules and in the limited understanding of their properties, which makes it difficult for domain experts to assess their viability (Lim et al., 2020).
@@
To address these challenges, another strategy for designing new molecules focuses on modifying/editing existing compounds using generative models (Bradshaw et al., 2019, Lim et al., 2020), reinforcement learning (Gottipati et al., 2020), genetic algorithms (Jensen, 2019), or from the domain experts themselves by trial and error (Bemis & Murcko, 1996, Schreiber, 2000, Welsch et al., 2010). These approaches are more likely to produce synthesizable molecules because the base molecule is known to exist and, therefore, can be more tractable for real-world applications. However, even with this strategy, a significant obstacle remains: sample-efficiency, i.e., how efficiently a method finds a promising molecular modification with a limited number of molecular property evaluations. Evaluating the properties of a chemical compound is a time-consuming and costly process, and many existing methods in the literature require numerous trials, making them less practical for accelerated discovery. Consequently, there is a critical need for a methodology that can generate target chemical compounds in a more sample-efficient manner.
@@
In this study, we introduce the Conditional Latent Space Molecular Scaffold Optimization (CLaSMO) method, a framework designed to address two critical challenges in chemical compound discovery: real-world applicability and sample-efficiency. CLaSMO combines a Conditional Variational Autoencoder (CVAE) (Higgins et al., 2016) with Latent Space Bayesian Optimization (LSBO) (Gómez-Bombarelli et al., 2018) to strategically modify input molecules and optimize their chemical properties.
@@
Figure 1: An example of the CLaSMO framework updating a scaffold for Quantitative Estimation of Drug-likeness (QED) optimization. CLaSMO identifies optimal regions in the CVAE latent space and selects bonding points on the scaffold. Chemical features from these points are used to guide the generation of substructures, which are then integrated into the scaffold (A, B, C, D) through small, targeted modifications to improve molecular properties.
@@
In drug discovery, a common strategy for generating synthesizable molecules is to work with molecular scaffolds—key substructures that serve as the foundation for chemical modifications and drug design (Bemis & Murcko, 1996). Building on this approach, our framework integrates small substructures into these scaffolds to improve key molecular properties. For modifications to be both effective and synthesizable, it’s crucial that the generative model understands how new substructures bond with the scaffold and enhance its properties. To achieve this, we condition substructure generation on the scaffold’s chemical features using a CVAE. Our novel data preparation and training strategy enables the CVAE to generate substructures that align with specific atomic environments. We further optimize this process using LSBO, which efficiently explores both the latent space of the CVAE and the scaffold’s chemical features. This allows CLaSMO to effectively select regions for modification and generate substructures that are chemically meaningful. By tracking molecular similarity between the initial scaffold and the optimized molecule, CLaSMO is designed to operate in low-budget scenarios and costly settings, such as wet-lab experiments, where high sample-efficiency is crucial. It ensures that modifications are both realistic and applicable to real-world contexts. Ultimately, this approach accelerates the discovery of novel, synthesizable compounds with improved properties. Figure 1 illustrates an example of molecular optimization with CLaSMO.
We evaluate our approach on a diverse suite of 20 molecular optimization tasks that span a wide range of objectives, including rediscovery of known compounds, multi-property optimization, and drug-likeness enhancement. These tasks present different challenges, enabling a comprehensive assessment of our model’s performance across varied molecular design scenarios. To further test robustness, we incorporate experimental settings with varying levels of similarity constraints between input and optimized molecules, reflecting both flexible and constrained optimization regimes. In addition to property optimization, we also assess the synthetic and retrosynthetic accessibility of the generated molecules. Our results show that CLaSMO consistently improves target properties in a sample-efficient manner, demonstrating strong generalizability and effectiveness across these diverse tasks.
@@ 2 Related Works
Molecular design strategies can broadly be divided into two categories: from-scratch-generation of molecules and modification-based approaches. Both categories have made significant strides in recent years, yet they also face unique challenges, particularly regarding real-world applicability and sample-efficiency.
@@
From-scratch-generation approaches focus on creating entirely new molecules by conducting a search to optimize the target property. A seminal work by Gómez-Bombarelli et al. (2018) introduced latent space optimization-based methodology, using a VAE to generate novel compounds by navigating the latent space of molecular representations. LSBO builds on this by efficiently reducing the number of expensive black-box evaluations required for molecular optimization, enabling the discovery of compounds with desirable properties in a continuous latent space. Since then, numerous studies have further refined and expanded the LSBO framework, focusing on method development and practical applications (Grosnit et al., 2021, Tripp et al., 2020, Maus et al., 2022, Griffiths & Hernández-Lobato, 2020, Boyar & Takeuchi, 2024, Boyar et al., 2024). However, like many other generation-from-scratch methods, such methodologies struggle with real-world applicability—i.e., the difficulty of synthesizing the generated molecules in real-world settings (Lim et al., 2020). Other generation methods, such as genetic algorithms (Jensen, 2019), Grammar VAE (Kusner et al., 2017), and Junction Tree (JT) VAE (Jin et al., 2018), aim to improve the chemical validity of generated structures. Recent advancements like GP-MOLFORMER (Ross et al., 2024) utilizes large language model-like approaches for molecular design, but the real-world applicability challenge remains a major limitation across these methodologies.
@@
Modification-based approaches, on the other hand, focus on adding substructures to existing molecules or scaffolds, often leading to more synthesizable and interpretable designs. Methods like Scaffold-GGM (Lim et al., 2020) employ a graph generative model to modify molecular scaffolds, thus improving properties with a higher chance of obtaining synthesizable molecules. Weller & Rohs (2024) introduces DrugHIVE, a deep hierarchical variational autoencoder that leverages scaffold modification to generate novel molecular compounds. A concurrent study by Lee et al. (2025) proposes a diffusion model based approach for both from-scratch and scaffold modification-based generation problems. There are many other scaffold-based optimization methodologies (Li et al., 2019, Langevin et al., 2020), not limited to generative modeling (Schreiber, 2000, Welsch et al., 2010, Miao et al., 2011). Techniques like Bradshaw et al. (2019)’s model ensure chemical validity in molecular modifications, and Gottipati et al. (2020)’s PGFS model uses reinforcement learning to guide additions to the base molecule. These approaches aim to avoid the low real-world viability problem faced by generation-from-scratch methods, as they build upon known molecular scaffolds, however, they lack advanced optimization methodologies that take sample-efficiency into account.
@@
CLaSMO combines the strengths of both categories, leveraging LSBO in the latent space of a CVAE for improved sample-efficiency while focusing on scaffold-based modifications. Unlike current LSBO-based molecular design methodologies in the literature, CLaSMO does not generate molecules from scratch but instead optimizes molecular properties by adding substructures to existing scaffolds. This approach mitigates the real-world applicability problem, and increases the chance of obtaining molecules that are both effective and practical for real-world synthesis. To evaluate the sample-efficiency of CLaSMO and benchmark it against established methodologies, we follow the sample-efficiency benchmark introduced by Gao et al. (2022). This benchmark provides a suite of challenging molecular property optimization tasks, using a variety of oracle functions, and compares the performance of numerous molecular design methods from the literature.
A key group of methods included in this benchmark are genetic algorithm-based approaches, which remain competitive and widely applicable to scaffold optimization. Notable examples include Graph-GA (Jensen, 2019), which uses graph-based genetic algorithms in combination with a generative model; Smiles-GA (Brown et al., 2019), which operates on SMILES string representations (Weininger, 1988); and STONED (Nigam et al., 2021), which leverages SELFIES (Krenn et al., 2020), a robust alternative to SMILES for molecular encoding. These methods have demonstrated strong performance and robustness, and benchmarking against them is highly recommended, as emphasized in Tripp & Hernández-Lobato (2023).
Beyond genetic algorithms, other scaffold-optimization-capable methodologies have also gained attention. For example, MARS Xie et al. (2021) employs a Markov Chain Monte Carlo (MCMC) approach with an annealing scheme and adaptive proposals to iteratively edit molecular graph fragments, guided by a graph neural network. It has demonstrated strong performance across various tasks, particularly in multi-objective optimization scenarios where properties such as bioactivity, drug-likeness, and synthesizability must be jointly optimized. Additionally, reinforcement learning techniques such as MolDQN (Zhou et al., 2019b) have proven effective for property optimization tasks. Recent advances in GFlowNets (Bengio et al., 2021) have also opened new avenues in molecular design. For instance, SynFlowNet (Cretu et al., 2024), a GFlowNet-based framework, supports scaffold optimization and has demonstrated competitive performance.
We adopt and modify the experimental setting proposed by Gao et al. (2022) in Section 5.2 to better align with the scope of our work—namely, molecular optimization in low-budget scenarios, where oracle evaluations are limited and sample-efficiency is essential. Within this setting, we demonstrate that CLaSMO consistently outperforms baseline methodologies across a diverse set of tasks. By combining a targeted scaffold-modification strategy with efficient search, CLaSMO offers a balanced and effective solution to molecular optimization under constrained evaluation budgets.
To assess the synthesizability of the generated molecules, we report both the Synthetic Accessibility score (SA score) Ertl & Schuffenhauer (2009) and the Retrosynthetic Accessibility score (RA score)(Thakkar et al., 2021). The SA score is obtained via estimation of synthetic feasibility based on the occurrence of molecular fragments in publicly available databases, serving as a fast and interpretable proxy for synthetic feasibility. However, it may underestimate difficulty for molecules with rare or unconventional motifs, or overestimate it for simple but impractical structures. To address such limitations, we complement this metric with the RA score, which reflects the predicted success of retrosynthetic route planning based on a classifier trained on reaction network outcomes. By incorporating both metrics, we obtain a more comprehensive assessment: SA scores reflect the structural simplicity and prevalence of a molecule in known chemical space, while RA scores provide insight into practical synthesis routes using available chemistry knowledge. This dual evaluation is particularly critical in low-data settings, where optimization strategies must not only find high-performing molecules, but also prioritize those amenable to real-world synthesis.
@@ 3.1 Conditional Variational Autoencoders (CVAEs)
A VAE (Kingma & Welling, 2014) consists of an encoder $f_{\phi}^{\text{enc}}:\mathcal{X}\to\mathcal{Z}$ and a decoder $f_{\theta}^{\text{dec}}:\mathcal{Z}\to\mathcal{X}$, where $\mathcal{X}$ represents the input space and $\mathcal{Z}$ the latent space. CVAEs extend the framework of VAEs by incorporating additional condition vector $\bm{c}$ into the latent variable model, facilitating the controlled generation of new instances. In the CVAE architecture, the encoder $q_{\phi}(\bm{z}|\bm{x},\bm{c})$ maps an input $\bm{x}$ and a condition $\bm{c}$ to a latent representation $\bm{z}$. Simultaneously, the decoder $p_{\theta}(\bm{x}|\bm{z},\bm{c})$ reconstructs $\bm{x}$ using both $\bm{z}$ and $\bm{c}$. The training of CVAEs is formulated as the minimization of the conditional variational lower bound:
$\textbf{J}(\theta,\phi;\bm{x},\bm{c})=-\mathbb{E}_{q_{\phi}(\bm{z}|\bm{x},\bm{c})}\left[\log p_{\theta}(\bm{x}|\bm{z},\bm{c})\right]+\mathrm{KL}\left(q_{\phi}(\bm{z}|\bm{x},\bm{c})\parallel p(\bm{z}\right)),$ (1)
@@ 3.2 Latent Space Bayesian Optimization (LSBO)
In BO, we start with numerous unlabeled instances $\{{\bm{x}_{i}}\}_{i\in[\mathcal{U}]}$ and a smaller set of labeled instances ${(\bm{x}_{i},y_{i})}_{i\in[\mathcal{L}]}$, where an input $\bm{x}_{i}\in\mathcal{X}$ represents a chemical compound, and a label $y_{i}\in\mathcal{Y}\subseteq\mathcal{R}$ indicates its properties such as docking scores. BO seeks to optimize a costly black-box function $f^{\rm BB}:\mathcal{X}\to\mathcal{Y}$, which corresponds to obtaining the physical properties of chemical compounds through experiments or time-consuming simulations in the context of molecular design problems. The goal is to maximize $f^{\rm BB}$ with minimal evaluations, using typically a Gaussian Process (GP) surrogate, trained using the labeled instances $\mathcal{L}$, to predict the function over $\mathcal{X}$. BO uses the surrogate to select an input $\bm{x}$ that may yield values surpassing the current maximum $\max_{i\in\mathcal{L}}y_{i}$. However, building a GP surrogate in high-dimensional spaces like chemical compounds is challenging. LSBO tackles this by employing a VAE/CVAE trained on the unlabelled instances $\mathcal{U}$ to reduce dimensionality, encoding instance in $\mathcal{X}$ to lower dimensional latent space $\mathcal{Z}$. This simplifies surrogate modeling and optimization because $\mathcal{Z}$ has lower-dimension than $\mathcal{X}$. During LSBO iterations, the acquisition function applied to the GP’s predictions selects new points in $\mathcal{Z}$ to evaluate. The chosen latent variable $\bm{z}_{i’}$ is decoded into a new input $\bm{x}_{i’}=f_{\theta}^{\rm dec}(\bm{z}_{i^{\prime}})$. This new instance is evaluated by $f^{\rm BB}$, and the results update $\mathcal{L}$ and refine the GP model. This cycle repeats until optimal results are achieved or resources are exhausted. In contexts like molecular design, LSBO aims to discover chemical compounds with optimal properties by efficiently navigating the reduced latent space.